What is the mass of
        
       (c) 10 moles of sodium sulphite (Na2SO3)?

Grade:12th pass

1 Answers

Harshit Singh
2 years ago
Dear Student


10 moles of sodium sulphite (Na2SO3)
Mass of 1 mole of Na2SO3= 2 × 23 + 32 + 3 ×16
=46 + 32 + 48 = 126 g
Mass of 10 moles of Na2SO3= 126 × 10
= 1260 g



Thanks